OpenStudy (anonymous):

I'm a bit confused on this one. Multiplication Property of Equality If 6x divided by 8 = 40, then _____. Choices: 6x times 8 = 320 40 = 6x times 8 6x = 320 40 = 6x divided by 8 Please explain. I think it is 6x = 320.

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

Thanks Hartnn! Could you please give me a brief explanation of the multiplication property of equality real quick? I just chose that because it made sense to me.

hartnn (hartnn):

yes,you can multiply a non zero number on both sides, here you multiplies by 8

hartnn (hartnn):

*multiplied

hartnn (hartnn):

6x/8 * 8 = 40 * 8 6x = 40*8 = 620

hartnn (hartnn):

*320

OpenStudy (anonymous):

Would you mind confirming another question please? I got no responses before.

OpenStudy (unklerhaukus):

\[\frac{6x} 8 = 40\]multiplying both sides by 8\[\frac{6x} 8\times8=40\times8\]canceling and factoring \[\frac{6x} {\cancel{8}}\times\cancel8=4\times8\times10\]\[6x=32\times10\]\[6x=320\]

hartnn (hartnn):

many are here to help u....

OpenStudy (anonymous):

Thanks Unkle!

OpenStudy (phi):

This is a definition question http://www.mathwords.com/m/multiplicative_property_of_equality.htm

OpenStudy (anonymous):

I know, but a lot of people would just come and not answer. I'm not saying that you have to answer it; only if you can and want to.

OpenStudy (anonymous):

Name the Property of Equality that justifies this statement: If p = q then p + s = q + s oReflexive Property oMultiplication Property oSymmetric Property oAddition Property I think it is the Addition Property. Thank you for the links!

OpenStudy (unklerhaukus):

yep

hartnn (hartnn):

because u added s on both sides

OpenStudy (anonymous):

Okay, thank you both very much! I greatly appreciate it!

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

@hartnn: Cool icon by the way. I like DBZ too. ^^

OpenStudy (unklerhaukus):

multiply both sides by a number \((8)\) or adding \((s)\) on both sides can be demonstrated with a pair of scales 

OpenStudy (unklerhaukus):